OpenStudy (anonymous):

Why does the multiplication of two binomials in the form (a + b) and (a - b) always result in a binomial?

OpenStudy (whpalmer4):

\[(a-b)(a+b) = a(a+b)-b(a+b) = a^2 + ab - ab - b^2\]The middle two terms always combine to 0, leaving just the two squares.